OpenStudy (anonymous):

If a problem states that x is over 3 and you need to get x alone, you multiply 3 to get x alone, but I'm not sure if it would be x or 3x

OpenStudy (anonymous):

\[\frac{ x }{ 3 } \times 3 = \frac{ 3x }{ 3 }\]cancel 3 to get x

OpenStudy (anonymous):

okay, thank you